OpenStudy (freemap):

The function n(d) = 4 • 3d – 1 represents the number of people who receive a newspaper on day d in a delivery schedule. What does the value 3 represent in this situation? Deliveries were made for 3 days. There are 3 people making deliveries. Three people received a newspaper on the first day in the schedule. The number of people receiving a newspaper on a given day is 3 times what it was the previous day.

OpenStudy (freemap):

I think A because day d in a delivery schedule

Directrix (directrix):

On day one of the delivery schedule, 4*3*1 -1 people received a paper. That would be 11 people on the first delivery schedule day. So, what would the 3 mean in front of the d?

OpenStudy (freemap):

D

Directrix (directrix):

If d = 2, 23 people get a newspaper on day 2 of the delivery schedule.

Directrix (directrix):

So, we have ruled out: Three people received a newspaper on the first day in the schedule.

Directrix (directrix):

The number of people receiving a newspaper on a given day is 3 times what it was the previous day. Day 1: 11 papers Day2: 23 Is 23 = 3 * 11 ?

OpenStudy (freemap):

no

OpenStudy (freemap):

23=3 *11= 33 not 23 so no not a true statment

OpenStudy (freemap):

Are you still there?

Directrix (directrix):

Yes. Would you look and see if the equation is this; n(d) = 4 • (3d – 1) I'm just wondering.

OpenStudy (freemap):

no, its not

Directrix (directrix):

I don't see any reason to think that either of these is corrrect. Deliveries were made for 3 days. There are 3 people making deliveries.

OpenStudy (freemap):

wait The function n(d) =4 •\[3^d-^1\]

OpenStudy (freemap):

i tried to make it where its 4 times 3d -1 but d and 1 up top not on bottom. If that makes since

Directrix (directrix):